OpenStudy (anonymous):

find the value of m that makes the expression a perfect square x2(2 is supposed to be a square)+4x+m

OpenStudy (anonymous):

x^2+4x+4 m = 4

OpenStudy (anonymous):

m=4

OpenStudy (anonymous):

\[(x+2)^2=x^2+4x+4\]